A box of mass m = 1.0 kg is pushed a distance d = 21 cm up an incline (angle θ = 17°) by a constant horizontal force P = 425 N. The coefficent of kinetic friction between the box and the incline is 0.24. Calculate the work done on the box by the applied force P.

Calculate the work done on the box by the frictional force.

Calculate the work done on the box by the gravitational force.


1
Expert's answer
2021-10-26T17:19:18-0400
"W_P=425(21)\cos{17}=8535\ J\\W_G=-(1)(9.8)(21)\sin{17}=-60\ J\\W_F=-(0.24)(1)(9.8)(21)=-2972\ J"